# FunDiff: Diffusion Models over Function Spaces for Physics-Informed Generative Modeling

### v1

@@ C.1 Baselines
C.2 Kolmogorov flow
@@
Figure 7: Kolmogorov flow. Unconditional generation without enforced divergence-free constraint. Examples of three generated velocity fields and their corresponding divergence fields.
Figure 8: Kolmogorov flow. Examples of the model predictions for three sample solutions. Each row shows a different sample, displaying from left to right: downsampled input velocity components ($u$, $v$), reference high-resolution velocity fields, predicted velocity fields, and the predicted divergence field. The model successfully reconstructs sharp flow features from the coarse input while strictly maintaining the divergence-free condition, as evidenced by the negligible values in the divergence fields (rightmost column).
Figure 9: Kolmogorov flow. Examples of the model predictions for three sample solutions. Each row shows a different sample with: downsampled input velocity components ($u$, $v$), reference high-resolution fields, predicted velocity fields, and the predicted divergence. While the model achieves basic upsampling of the velocity fields, the reconstructions exhibit noticeable blurring compared to the sharp features in the reference solutions. The large values in the divergence fields (rightmost column) indicate that the incompressibility constraint is not enforced.
C.3 Burgers’ equation
@@
Following Wang et al. [65], we generate our dataset by sampling 4,000 initial conditions from a GRF with covariance operator $\mathcal{N}(0,25^{2}(-\Delta+5^{2}I)^{-4})$. For each initial condition, we solve the Burgers’ equation in Eq. (4.2) using spectral methods with periodic boundary conditions. The numerical solution is computed using the Chebfun package [66] with a spectral Fourier discretization and the fourth-order ETDRK4 time-stepping scheme [67], using a time-step size of $10^{-4}$. We record solution snapshots every $\Delta t=0.005$ up to $t=1$, resulting in 201 temporal snapshots for each realization on a uniform $200\times 201$ spatio-temporal grid.
Figure 10: Burgers’ equation. Representative generated PDE solutions and the corresponding PDE residuals, given the coarse measurements in $50\times 50$ space-time resolution. The model is trained with incorporated PDE residual constraints.
Figure 11: Burgers’ equation. Representative generated PDE solutions and the corresponding PDE residuals, given the coarse measurements in $50\times 50$ space-time resolution. The model is trained without incorporating PDE residual constraints.
C.4 Linear elasticity
@@
$E(\mathbf{x}),\nu(\mathbf{x})=\begin{cases}E_{f},\nu_{f},&\text{ if }\mathbf{x}\in\Omega_{f},\\ E_{m},\nu_{m},&\text{ if }\mathbf{x}\in\Omega_{m},\end{cases}$
where ${\Omega}_{f}$ and ${\Omega}_{m}$ represent fiber and matrix phases, respectively. Our dataset comprises 7,500 unique RVEs, each discretized on a uniform $256\times 256$ grid. For all RVEs, the matrix material properties are fixed with $E_{m}=1$ (GPa) and $\nu_{m}=\nu_{f}=0.35$. The volume fraction $vof$ and the Young’s modulus $E_{f}$ of fiber are randomly generated through the Latin hypercube sampling (LHS) design method, ensuring that $vof\in[0.1,0.65]$ and $E_{f}\in[10^{-6},10^{6}]$. The $A_{ijkl}(\mathbf{x})$ of each RVE is computed by solving Eq. (4.4) subjected to Eq. (4.5) using a Fast Fourier Transform (FFT)-based homogenization method [62]. A detailed and comprehensive implementation of generating $A_{ijkl}(\mathbf{x})$ can be found in Ref. [68]. Given that the diagonal components $A_{11}$, $A_{22}$, and $A_{33}$ of the strain concentration tensor dominate the mechanical response, these components are exclusively used to train our functional autoencoder.
Figure 12: Linear elasticity. Generative samples obtained by training a discrete ViT autoencoder and latent diffusion model using the pipeline in Fig. 1. Noticeable discontinuities appear in the generated outputs, manifesting as distinct patch artifacts.
C.5 Turbulence mass transfer

### v2

@@ Acknowledgments
Continuous Vision Transformer (CViT) [61].
A neural field architecture that integrates a ViT encoder with trainable grid and coordinate embeddings for continuous, resolution-agnostic querying. In our implementation, the encoder operates on $16\times 16$ patches with an embedding dimension of 256, depth 8, and 8 attention heads (MLP ratio 2). The decoder adopts a similar configuration with an embedding dimension of 256, depth 4, and 8 attention heads, followed by two MLP layers to obtain the final output.
CNextU-Net [63].
A modernized U-Net backbone built upon ConvNeXt blocks [83], providing a high-capacity convolutional baseline competitive with transformer-based architectures for spatiotemporal field regression. our model employs a four-stage encoder–decoder hierarchy with two ConvNeXt blocks per stage and two additional blocks at the bottleneck, initialized with 32 feature channels and doubling features at each downsampling level.
AviT [64].
An attention-based ViT variant that sequentially attends over spatial and temporal dimensions, enabling efficient modeling of dynamic physical systems. Our implementation follows the base model described in Ref. [64], consisting of 8 AViT blocks with hidden dimension 512, 8 attention heads, and grouped axial attention across spatial partitions.
DPOT [62].
A PDE foundation operator trained via autoregressive denoising and Fourier attention, designed to generalize across diverse PDE systems and boundary conditions. In our implementation, DPOT processes inputs of size $256\times 256$ with a patch size of 16, embedding dimension 512 and an MLP ratio of 2. The core architecture consists of a Fourier-attention transformer stack with 8 layers, interleaved with 4 Fourier mixer blocks (via AFNO) for spatial-frequency mixing.
Diffusion Posterior Sampling (DPS) [65].
An inverse solver for general operators $\mathcal{A}$ that augments the reverse diffusion process with a likelihood-score guidance term $\nabla_{x}\log p_{\sigma}(y\,|\,\mathcal{A}(x))$ to approximate posterior sampling. We adopt the original denoising diffusion probabilistic model (DDPM) backbone with four encoder blocks and feature depths $(32,64,128,256)$, each comprising two residual blocks and group normalization (8 groups). Attention is applied in the last two stages with eight heads, and the model uses an embedding width of 256 and one middle residual block. Sampling is performed over 1,000 diffusion steps using a cosine noise schedule and a tuned guidance scale of $\lambda=1000$, selected for best reconstruction accuracy.
DiffusionPDE [42].
A conditional diffusion model that jointly learns the distribution of PDE coefficients and solutions, guided by both observation consistency and PDE residual constraints during denoising. We follow the official implementation with a four-stage U-Net backbone using feature channels $(64,128,256,512)$, two residual blocks per stage, and self-attention in the bottleneck layer. Training is conducted for 1,000 diffusion steps with a cosine noise schedule and dual guidance weights tuned to balance data fidelity and physics consistency, selected for optimal reconstruction accuracy.
Cross-Attention Diffusion (Cross-Attn) [51].
A diffusion framework that employs cross-attention to reconstruct spatially distributed physical fields from sparse or irregular measurements. Following the original implementation, we use the aforementioned U-Net as the backbone. For the conditioning, measurement tokens are obtained through patch embeddings and processed by 4 transformer layers with an embedding dimension of 256, MLP ratio 2, and 8 attention heads. Cross-attention modules are integrated into each U-Net residual block to encode measurement-conditioned context and guide the denoising process.
S4.2 Kolmogorov flow
@@
Figure S1: Kolmogorov flow. Unconditional generation without enforced divergence-free constraint. Examples of three generated velocity fields and their corresponding divergence fields.
Figure S2: Kolmogorov flow. Examples of the model predictions for three sample solutions. Each row shows a different sample, displaying from left to right: downsampled input velocity components ($u$, $v$), reference high-resolution velocity fields, predicted velocity fields, and the predicted divergence field. The model successfully reconstructs sharp flow features from the coarse input while strictly maintaining the divergence-free condition, as evidenced by the negligible values in the divergence fields (rightmost column).
Figure S3: Kolmogorov flow. Examples of the model predictions for three sample solutions. Each row shows a different sample with: downsampled input velocity components ($u$, $v$), reference high-resolution fields, predicted velocity fields, and the predicted divergence. While the model achieves basic upsampling of the velocity fields, the reconstructions exhibit noticeable blurring compared to the sharp features in the reference solutions. The large values in the divergence fields (rightmost column) indicate that the incompressibility constraint is not enforced.
S4.3 Burgers’ equation
@@
Following Wang et al. [84], we generate our dataset by sampling 4,000 initial conditions from a GRF with covariance operator $\mathcal{N}(0,25^{2}(-\Delta+5^{2}I)^{-4})$. For each initial condition, we solve the Burgers’ equation in (5) using spectral methods with periodic boundary conditions. The numerical solution is computed using the Chebfun package [85] with a spectral Fourier discretization and the fourth-order ETDRK4 time-stepping scheme [86], using a time-step size of $10^{-4}$. We record solution snapshots every $\Delta t=0.005$ up to $t=1$, resulting in 201 temporal snapshots for each realization on a uniform $200\times 201$ spatio-temporal grid.
Figure S4: Burgers’ equation. Representative generated PDE solutions and the corresponding PDE residuals, given the coarse measurements in $50\times 50$ space-time resolution. The model is trained with incorporated PDE residual constraints.
Figure S5: Burgers’ equation. Representative generated PDE solutions and the corresponding PDE residuals, given the coarse measurements in $50\times 50$ space-time resolution. The model is trained without incorporating PDE residual constraints.
S4.4 Linear elasticity
@@
[…]
where ${\Omega}_{f}$ and ${\Omega}_{m}$ represent fiber and matrix phases, respectively. Our dataset comprises 7,500 unique RVEs, each discretized on a uniform $256\times 256$ grid. For all RVEs, the matrix material properties are fixed with $E_{m}=1$ (GPa) and $\nu_{m}=\nu_{f}=0.35$. The volume fraction vofvof and the Young’s modulus $E_{f}$ of fiber are randomly generated through the Latin hypercube sampling (LHS) design method, ensuring that $vof\in[0.1,0.65]$ and $E_{f}\in[10^{-6},10^{6}]$. The $A_{ijkl}(\mathbf{x})$ of each RVE is computed by solving (7) subjected to (8) using a Fast Fourier Transform (FFT)-based homogenization method [67]. A detailed and comprehensive implementation of generating $A_{ijkl}(\mathbf{x})$ can be found in Ref. [87]. Given that the diagonal components $A_{11}$, $A_{22}$, and $A_{33}$ of the strain concentration tensor dominate the mechanical response, these components are exclusively used to train our functional autoencoder.
Figure S6: Linear elasticity. Generative samples obtained by training a discrete ViT autoencoder and latent diffusion model using the pipeline in Fig. 1. Noticeable discontinuities appear in the generated outputs, manifesting as distinct patch artifacts.
S4.5 Turbulence mass transfer